\begin{thebibliography}{999999}

\markboth{}{}



\bibitem[E]{Escobar}
J. F. Escobar, Uniqueness theorems on conformal deformation of metrics, Sobolev inequalities, and an eigenvalue estimate, Comm. Pure Appl. Math., {\bf{43}}  (7) (1990) 857-883.

\bibitem[O]{O} 
M. Obata, Certain conditions for a manifold to be isometric with a sphere, J. Math. Soc. Japan, {\bf{14}} (1962) 333-340.


\bibitem[R1]{Re1}
R. C. Reilly, Applications of the Hessian operator in a Riemannian manifold, Indiana University Mathematics Journal, {\bf{26}} (3) (1977) 459-72.

\bibitem[R2]{Re2}
R. C. Reilly, Geometric applications of the solvability of Neumann problems on a Riemannian manifold, Archive for Rational Mechanics and Analysis, {\bf{75}} (1)  (1980) 23-29.

\bibitem[X]{Xia}
Changyu Xia, The first nonzero eigenvalue for manifolds with Ricci curvature having positive lower bound, Chinese Math. into the 21st Century, 1991, Tian Jing. Proceedings of the International Conference on Chinese Math. into the 21st Century, 1988, 237-243.

\bibitem[X-H]{XinHongwei}
Ren Xin'an and Xu Hongwei, A lower bound for first eigenvalue with mixed boundary condition, Appl. Math. J. Chinese Univ., Ser. B, {\bf{19}} (2) (2004) 223-228.



\end{thebibliography}

\documentclass{amsart}
\usepackage[english]{babel}
\usepackage{amsfonts, amsmath, amsthm, amssymb,amscd,indentfirst}

\usepackage{esint}
\usepackage{graphicx}
\usepackage{epstopdf}
\usepackage{amsmath,amssymb,latexsym,indentfirst}
\usepackage{times}
\usepackage{palatino}
\usepackage[colorlinks]{hyperref}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}[theorem]{Definition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{remark}[theorem]{Remark}

\def\d{\partial}

\def\bp{\begin{proof}}
\def\ep{\end{proof}}

\def\d{\partial}

\begin{document}

\title[Rigidity on an eigenvalue problem with mixed boundary condition]{Rigidity on an eigenvalue problem with mixed boundary condition}

\author{\textsc{S\'{e}rgio Almaraz}}
\address{Universidade Federal Fluminense (UFF), Instituto de Matem\'{a}tica, Campus do Gragoat\'a\\
              Rua Prof. Marcos Waldemar de Freitas S/N ,  24210-201, Niter\'{o}i-RJ, Brazil.}
              \email{sergio.m.almaraz@gmail.com}
\author{\textsc{Ezequiel Barbosa}}
\address{Universidade Federal de Minas Gerais (UFMG), Departamento de Matem\'{a}tica, Caixa Postal 702, 30123-970, Belo Horizonte-MG, Brazil.}
\email{ezequiel@mat.ufmg.br}
\thanks{The two authors were partially supported by  CNPq/Brazil grants.}

\begin{abstract}
We prove an Obata-type rigidity result for the spherical cap and apply it for an eigenvalue problem with mixed boundary condition.

MSC classes: 53C24, 35P15.
\end{abstract}

\maketitle

\section{Introduction and statements of the results}\label{int}

Let $(M^n,g)$ be a compact Riemannian manifold, with a possibly non-empty boundary $\Sigma$.  We denote by $\text{Ric}_g$ and $R_g$ respectively the Ricci and scalar curvatures of $(M, g)$. When $\Sigma\neq \emptyset$, we also assume that $\Sigma$ is oriented by an outward pointing unit normal vector $\nu$, so that  its mean curvature is $H_{\Sigma}=\text{div}_g\nu$.

The classical Obata's rigidity theorem is a very well-known result with many important applications:

\begin{theorem}[\cite{O}]\label{thm:obata}
A connected closed Riemannian manifold $(M^n,g)$ is isometric to the standard unit sphere $(\mathbb{S}^n,g_0)$ if and only if there is a non-constant function $\phi$ satisfying
\[
\nabla^2 \phi+\phi g=0,
\]
where $\nabla^2$ is the Hessian with respect to the metric $g$.
\end{theorem}
\begin{corollary}
If a connected closed Riemannian manifold $(M^n,g)$ satisfies 
$$\text{Ric}_g\geq (n-1)g,$$ 
then the first eigenvalue $\lambda_1(g)$ of the Laplacian $\Delta_g$ satisfies
$\lambda_1(g)\geq n$, and equality holds if and only if $(M,g)$ is isometric to $(\mathbb{S}^n,g_0)$.
\end{corollary}

To the best of our knowledge, there are two versions of those results for manifolds with non-empty boundary: 

\begin{theorem}[\cite{Re1, Re2}]
A connected compact Riemannian manifold $(M^n,g)$ with non-empty boundary $\Sigma$  is isometric to the standard unit hemisphere $(\mathbb{S}^n_+,g_0)$ if and only if there is a non-constant function $\phi$ satisfying
\[
\begin{cases}
\nabla^2 \phi+\phi g=0\:\:\:\text{in}\:M,
\\
\phi= \text{const.\:\:\:on\:} \Sigma.
\end{cases}
\]
\end{theorem}
\begin{corollary}
Let $(M^n,g)$ be a connected compact Riemannian manifold with nonempty boundary $\Sigma$. Assume that $\text{Ric}_g\geq(n-1)g$ and $H_{\Sigma}\geq 0$. Then the first eigenvalue $\lambda_1(g)$ of $\Delta_g$ with Dirichlet boundary condition satisfies $\lambda_1(g)\geq n$. Moreover, $\lambda_1(g)=n$ if and only if $(M,g)$ is isometric to $(\mathbb{S}^n_+,g_0)$.
\end{corollary}
\begin{theorem}[\cite{Escobar, Xia}]
A connected compact Riemannian manifold $(M^n,g)$ with non-empty boundary  $\Sigma$ is isometric to  $(\mathbb{S}^n_+,g_0)$ if and only if \[
\begin{cases}
\nabla^2 \phi+\phi g=0\:\:\:\text{in}\:M,
\\
\frac{\partial}{\partial \nu}\phi=0\text{\:\:\:on\:} \Sigma.
\end{cases}
\]
\end{theorem}
\begin{corollary}
Let $(M^n,g)$ be a connected compact Riemannian manifold with nonempty boundary $\Sigma$. Assume that $\text{Ric}_g\geq(n-1)g$ and $H_{\Sigma}\geq 0$. Then the first eigenvalue $\lambda_1(g)$ of $\Delta_g$ with Neumann boundary condition  satisfies $\lambda_1(g)\geq n$. Moreover $\lambda_1(g)=n$ if and only if $(M,g)$ is isometric to $(\mathbb{S}^n_+,g_0)$.
\end{corollary}

Our main result extends the above to a mixed boundary condition, under certain hypothesis on the boundary mean curvature:

\begin{theorem}\label{main:thm}
Let $\theta\in (0,\frac{\pi}{2})$. A connected compact Riemannian manifold $(M^n,g)$ with non-empty $\Sigma$ satifying $H_{\Sigma}\geq (n-1)\tan(\theta)$ is isometric to a geodesic ball in $(\mathbb{S}^n,g_0)$, whose boundary is a totally umbilical hypersphere with mean curvature $(n-1)\tan(\theta)$, if and only if there is a non-constant function $\phi$ satisfying
\[
\begin{cases}
\nabla^2 \phi+\phi g=0\:\:\:\text{in}\:M,
\\
\tan(\theta)\frac{\partial}{\partial \nu}\phi+\phi=0\text{\:\:\:on\:} \Sigma.
\end{cases}
\]
\end{theorem}
As an application, we are able to complete the study initiated by Xin'an and Hongwei \cite{XinHongwei} on the first eigenvalue of the Laplacian operator with mixed boundary condition. Combining the inequality proved in \cite{XinHongwei} with our rigidity statement, we obtain

\begin{theorem}\label{main:application}
Let $(M^n,g)$ be a connected compact Riemannian manifold with nonempty boundary $\Sigma$ whose 2nd fundamental form and mean curvature are denoted by $h_{\Sigma}$ and $H_{\Sigma}=\text{tr}\,h_{\Sigma}$ respectively. If for some $\theta\in (0,\frac{\pi}{2})$,
$$\text{Ric}_g\geq (n-1)g,  \:\:\:h_{\Sigma}\geq-2\cot(\theta)g \:\:\text{and}\:\: H_{\Sigma}\geq(n-1)\tan(\theta),$$ 
then the first eigenvalue $\lambda_1(g,\theta)$ of the problem
\begin{equation}\label{eigenvalue:problem}
\begin{cases}
\Delta_g\phi+\lambda\phi=0&\text{in}\:M,
\\
\tan(\theta)\frac{\partial}{\partial\nu}\phi+\phi=0&\text{on}\:\Sigma,
\end{cases}
\end{equation}
satisfies $\lambda_1(g,\theta)\geq n$. Moreover, the equality $\lambda_1(g,\theta)=n$ holds if and only if $M$ is isometric to an $n$-dimensional spherical cap of $(\mathbb{S}^n, g_0)$, whose boundary is a totally umbilical hypersphere with mean curvature $(n-1)\tan(\theta)$.
\end{theorem}

Our contribution is the equality part, since the inequality was proved in \cite{XinHongwei}. 

In Section 2 we prove Theorem \ref{main:thm} and in Section 3 we prove the rigidity part of Theorem \ref{main:application}


\section{Proof of Theorem \ref{main:thm}}
Without loss of generality, we assume that $\phi(p)=\max\limits_{M}\phi=1$ for some $p\in M$. Since $\tan(\theta)>0$, it follows from  the second equation in \eqref{eigenvalue:problem} that $p\notin\Sigma$. 

Let $\gamma:[0,l_0]\rightarrow M$ be a geodesic such that $|\gamma'|=1$, $\gamma(0)=p$. Define $u(t)=\phi\circ \gamma(t)$. Since $\nabla^2 \phi+\phi g=0$, we obtain the o.d.e.
\begin{equation*}
\begin{cases}
u''+u=0,
\\
u(0)=1, \:\:u'(0)=0.
\end{cases}
\end{equation*}
Hence, $u(t)=\cos t$ in $[0,l_0]$.
Since $M$ is compact, we can choose $\gamma$ in such a way that $l_0=dist_g(p,\Sigma)=dist_g(p,q)$, where $q=\gamma(l_0)\in \Sigma$. 

Set
\[
\overline{l}=\sup \{l\in (0,l_0]\,|\,\exp_{p}: B_l(p)\subset M\rightarrow B_l(0)\subset \mathbb{R}^n \,\,\mbox{is a diffeomorphism} \}\,.
\]
Observe that $\overline{l}>0$, because the exponential map $\exp_{p}: B_l(p)\subset M\rightarrow B_l(0)\subset \mathbb{R}^n$ is a diffeomorphism for small $l$. 
Now, taking any point $p_0\in \mathbb{S}^n$ and using the exponential map on $S^n$ at $p_0$, we obtain a diffeomorphism  
$$
\psi: B_l(p)\subset M\rightarrow B_l(p_0)\subset \mathbb{S}^n$$ 
for any $0<l<\overline{l}$ with $l<\pi$. As in the proof of the classical Obata's Theorem \cite[pp.338-339]{O}, we can prove that $\psi$ is an isometry. So $\overline{l}<\pi$, otherwise $M$ is closed which is a contradiction.

It follows from definition that $\overline{l}\leq l_0$. We claim that $\overline{l}=l_0$. Suppose by contradiction $\overline{l}< l_0$. Then there exists a geodesic starting at $p$ with a cut-locus point in $\bar  B_{\overline{l}}(p)$. Hence, we see that $\overline{l}\geq\pi$ which is a contradiction.
Thus, $\overline{l}=l_0$.

\vspace{0.1cm}
\emph{Claim:} We have $H_{\partial B_{l_0}(p)}=\tan(\theta)(n-1)$.

Since $\psi$ is an isometry, it is enough to prove that $\overline l=\pi/2-\theta$. Note that $u(t)=\cos(t)$ implies
\begin{equation}\label{eq:1}
-\tan(\theta)\frac{\partial}{\partial \nu}\phi(q)=-\tan(\theta)u'(\overline{l})=\tan(\theta)\sin (\overline{l})
\end{equation}
because $\gamma$ intersects $\Sigma$ orthogonally at $\gamma(\overline{l})$.
On the other hand, it holds
\begin{equation}\label{eq:2}
\phi(q)=u(\overline{l})=\cos(\overline{l})
\end{equation}
It follows from the second equation in \eqref{eigenvalue:problem} that the l.h.s. of \eqref{eq:1} and \eqref{eq:2} equal, so 
$$
\tan(\theta)=\cot(\overline{l})=\tan(\pi/2-\overline{l}).
$$
Now, because $\tan$ is an injective function in $(-\pi/2,\pi/2)$, we have $\overline{l}=\pi/2-\theta$.
This proves the claim.

We are now able to finish the proof of Theorem \ref{main:thm}. Using $\overline{l}=l_0=dist_g(p,q)$ we obtain $\Sigma \subset M\setminus B_{l_0}(p)$ and $q\in \partial B_{l_0}(p)\cap \Sigma$. Using $H_{\partial B_{l_0}(p)}=\tan(\theta)(n-1)\leq H_{\Sigma}$ and the maximum principle, we conclude that $\partial B_{l_0}(p)=\Sigma$ and, consequently, $M=\bar B_{l_0}(p)$ which is isometric to $\bar B_{l_0}(p_0)\subset \mathbb{S}^n$.



\section{The rigidity part in Xin'an-Hongwei's theorem}

\begin{theorem}[\cite{XinHongwei}]\label{thm:XinHongwei}
\footnote{
Although our hypothesis on $h_{\Sigma}$ are weaker than stated in \cite{XinHongwei} (which assumes $h_{\Sigma}\geq 0$), the proof of Theorem \ref{thm:XinHongwei} follows exactly the same lines as \cite{XinHongwei}.
}
Let $(M^n,g)$ be a connected compact Riemannian manifold with nonempty boundary $\Sigma$. If for some $\theta\in (0,\frac{\pi}{2})$
$$\text{Ric}_g\geq (n-1)g,  \:\:\:h_{\Sigma}\geq-2\cot(\theta)g \:\:\text{and}\:\: H_{\Sigma}\geq(n-1)\tan(\theta),$$ 
then the first eigenvalue of problem \eqref{eigenvalue:problem} satisfies 
$\lambda_1(g,\theta)\geq n$. 
\end{theorem}
\begin{proof}
As in \cite[p.227]{XinHongwei}, setting $u=\frac{\d f}{\d \nu}\Big|_{\Sigma}$ and $z=f|_{\Sigma}$, and using \cite{Re1}, we obtain
\begin{align*}
\int_M((\Delta f)^2-|\nabla^2f|^2)dv_g
=&\int_{\Sigma}\big((\bar\Delta z+H_{\Sigma}u)u-<\bar\nabla z,\bar\nabla u>+h_{\Sigma}(\bar\nabla z,\bar\nabla z)\big)d\sigma
\\
&+\int_M\text{Ric}_g(\nabla f,\nabla f)dv_g
\end{align*}
where the bars stand for the restrictions to $\Sigma$. If $f$ satisfies
$\Delta_gf+\lambda_1(g,\theta)f=0$ and $u+z\cot(\theta)=0\:\: \text{on}\:\Sigma$, using $\text{Ric}_g\geq (n-1)g$ we obtain
\begin{align*}
\frac{n-1}{n}\lambda_1(g,\theta)(\lambda_1(g,\theta)-n)\int_Mf^2dv_g
\geq&\int_{\Sigma}(2\cot(\theta)|\bar\nabla z|^2+h_{\Sigma}(\bar\nabla z,\bar\nabla z))d\sigma
\\
&+\int_{\Sigma}(H_{\Sigma}-(n-1)\tan(\theta))u^2d\sigma.
\end{align*}
From the hypotheses on $h_{\Sigma}$ and $H_{\Sigma}$ we conclude that 
$\lambda_1(g,\theta)(\lambda_1(g,\theta)-n)\geq0$.
\end{proof}
As a consequence of Theorem \ref{main:thm} we prove:
\begin{theorem}
Equality  $\lambda_1(g,\theta)=n$  in Theorem \ref{thm:XinHongwei} holds if and only if $M$ is isometric to an $n$-dimensional spherical cap of $\mathbb{S}^n$, whose boundary is a totally umbilical hypersphere with mean curvature $(n-1)\tan(\theta)$.
\end{theorem}
\begin{proof}
Note from the previous proof that the equality holds if and only if the following four conditions hold:

\begin{itemize}
\item[(i)] $\text{Ric}_g(\nabla f, \nabla f)=(n-1)|\nabla f|^2$;
\item[(ii)] $H_{\Sigma}=(n-1)\tan(\theta)$;
\item[(iii)] $2\cot(\theta)|\bar\nabla z|^2+h_{\Sigma}(\bar\nabla z,\bar\nabla z)=0$;
\item[(iv)] $|\nabla^2f|^2=\frac{(\Delta f)^2}{n}$.
\end{itemize}
We can see that (iv) is equivalent to
\[
\nabla^2f=\frac{\Delta f}{n}g\,.
\]
Then we apply Theorem \ref{main:thm} to get the result.
\end{proof}




\end{document}